OpenStudy (anonymous):

Which of the following is a solution of x2 + 2x + 8? 4 + i times the square root of 7 −1 + i times the square root of 7 1 + i times the square root of 7 −4 + i times the square root of 7

Nnesha (nnesha):

use the quadratic formula \[\huge\rm \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] \[\huge\rm Ax^2+Bx+C=0\] where a=leading coefficient b=middle term c=constant

Nnesha (nnesha):

and x^2 is same as 1x^2 there is invisible one